# Can Neural Decompilation Assist Vulnerability Prediction on Binary Code?

### v1

@@ 8. Conclusion
In this work, we presented an experimental study on vulnerability prediction in binary files. In particular, we leveraged neural decompilation to translate binary code into source code, treating both programs as simple text. Our goal was to demonstrate that a more effective solution for detecting vulnerabilities in binary files is to use neurally decompiled source code, rather than relying on complex toolchains to convert binary code into an intermediate representation or to extract control flows for decompilation. Our experimental analysis revealed that neural decompilation is fit for this task, achieving great results comparable with state-of-the-art works, and that transformer-based and DL models can accurately predict whether a binary file is vulnerable and identify the type of vulnerabilities, mapped to the MITRE CWE taxonomy, using the neurally decompiled code.

### v2

@@ 6. Conclusion
In this work, we presented an experimental study on vulnerability prediction in binary files. In particular, we leveraged neural decompilation to translate binary code into source code, treating both as simple text. Our goal was to demonstrate that a more effective solution for detecting vulnerabilities in binary files is to use neurally decompiled source code, rather than relying on complex toolchains to convert binary code into an intermediate representation or to extract control flows for decompilation. Our experimental analysis revealed that neural decompilation is fit for this task, achieving great results comparable with state-of-the-art works, and that transformer-based and DL models can accurately predict whether a binary file is vulnerable and identify the type of vulnerabilities, mapped to MITRE CWE, using the neurally decompiled code.
Acknowledgments
This work has been partially supported by the PNRR - DM 118/2023 project, Mis.: I.3.4 Dottorati Transizione Digitale (CUP E66E23001030002), and the GENIO project, funded by MIMIT (CUP B69J23005770005).